### v1

# AirIO: Learning Inertial Odometry with Enhanced IMU Feature Observability air-io.github.io

Abstract
Inertial odometry (IO) using only Inertial Measurement Units (IMUs) offers a lightweight and cost-effective solution for Unmanned Aerial Vehicle (UAV) applications, yet existing learning-based IO models often fail to generalize to UAVs due to the highly dynamic and non-linear-flight patterns that differ from pedestrian motion. In this work, we identify that the conventional practice of transforming raw IMU data to global coordinates undermines the observability of critical kinematic information in UAVs. By preserving the body-frame representation, our method achieves substantial performance improvements, with a 66.7% average increase in accuracy across three datasets. Furthermore, explicitly encoding attitude information into the motion network results in an additional 23.8% improvement over prior results. Combined with a data-driven IMU correction model (AirIMU) and an uncertainty-aware Extended Kalman Filter (EKF), our approach ensures robust state estimation under aggressive UAV maneuvers without relying on external sensors or control inputs. Notably, our method also demonstrates strong generalizability to unseen data not included in the training set, underscoring its potential for real-world UAV applications.
Index Terms:
Learning-based Inertial Odometry
I Introduction
Inertial Measurements Units (IMUs) are inexpensive and ubiquitous sensors that provide linear accelerations and angular velocities. Due to the size, weight, power, and cost (SWAP-C) constraints, IO based only on light and low-cost sensors is ideal for UAVs applications ranging from 3D mapping [1], exploration [2] to physical interaction [3]. Compared to exteroceptive sensors like vision and LiDAR, IMUs are unaffected by visual degradation factors such as motion blur [4] or dynamic object interruption [5], which are common in agile UAV flights [6]. Despite these advantages, current IO solutions struggle to accurately learn and adapt to the complex motion models of UAVs due to the inherent noise and bias of the IMU measurement, leading to large drift over time, particularly during agile maneuvers.
Figure 1: The upper figure shows quadrotor tilt dynamics: take-off, hovering, and landing maintain initial orientation; acceleration increases tilt for thrust; cruising holds steady tilt for constant velocity; and deceleration tilts oppositely to create drag. The lower figure depicts a simulated straight-line flight with constant yaw from the Pegasus Simulator [7]. The red dashed line tracks pitch angle adjustments, while black solid line shows velocity changes over time.
Most recent advances in learning-based IO have focused on pedestrian and legged robots [8, 9, 10], utilizing motion priors like stride length [11] and repetitive gait patterns [12, 13] beyond IMU pre-integration [14]. In contrast, multirotor UAV motion involves rigorous maneuvers affecting orientation and thrust, resulting in highly dynamic, nonlinear behavior without clear priors. Small attitude changes can cause significant variations in velocity and position, complicating the IO process and making pedestrian-focused methods less effective [6]. As a result, deploying learning-based IO on UAVs often requires additional sensors, such as tachometers [15] or control inputs like thrust commands [6]. This raises our main questions:
Why is learning IO not directly applicable to drones?
How can learning IO be effectively deployed for UAVs?
In this paper, we investigate the effective representation of learning-based IO in multirotor UAV motion and propose a solution that relies solely on IMU data. Through rigorous feature analysis and examination of UAV dynamics, we identify that the commonly used global-frame representation is less effective for dynamic, agile maneuvers due to the less observable attitude information compared to IMU data in the body-frame representation. To address this, we introduce AirIO, a learning-based IO method that explicitly encodes attitude information and predicts velocity using body-frame representation. We further integrate an EKF that fuses IMU pre-integration with the learning-based model for odometry estimation. We show that our approach outperforms state-of-the-art algorithms [6] in various scenarios and environments, even without additional sensors or control information.
The main contributions of this work are:
  1. 1.
    We conduct the analysis of the extracted features from different representations using t-SNE and principal component analysis (PCA). The results show that body-frame representation is more expressive and observable. Simply changing the input representation can significantly improve IO accuracy by an average of 66.7%.
  2. 2.
    Grounded on this effective representation, we develop a learning-based IO framework AirIO that explicitly encodes the attitude information and fuses it with the body-frame IMU data to predict the velocity. Explicitly encoding the attitude information further improves accuracy by an additional 23.8%.
  3. 3.
    We further integrate an uncertainty-aware IMU preintegration model and the learned motion network into EKF for odometry estimation. Extensive real-world experiments validate our system, which outperforms existing IO algorithms without the need for additional sensors or control information. Our model also demonstrates generalizability to the unseen datasets that are not included in the training sets.
II Related Works
II-A Model-based Inertial Odometry
Traditional model-based inertial odometry (IO) methods rely on kinematic motion models [14] to estimate relative motion and are often integrated with exteroceptive sensors such as cameras [16] and LiDAR [1]. While these algorithms achieve high accuracy, their dependence on external sensors makes them vulnerable to disturbances caused by agile motion or dynamic environments. To address the SWAP challenge of lightweight multirotor UAV navigation, Svacha et al. [17, 18] introduced a method that estimates tilt and velocity by fusing tachometer and IMU data. Although effective, this approach still rely on auxiliary sensor, limiting their applicability in environments where such sensors are unavailable or unreliable.
II-B Learning Inertial Odometry
Recent advances in deep learning have sparked research in data-driven methods for learning velocity and displacement from inertial data [11, 19, 9, 20]. Approaches such as IONet [19], A2DIO [21], and NILoc [20] formulate inertial navigation as a sequential problem task by predicting relative position displacements using IMU data. Subsequent work has incorporated these learned displacements into EEKF [8, 22, 23] or batched optimization frameworks [24]. To ensure consistency in IMU measurements, many of these methods transform the input IMU data into global coordinates. For pedestrian navigation, RoNIN [9] proposed a heading-agnostic coordinate frame (HACF), aligning the accelerometer’s gravity vector with the z-axis and constraining rotation to the horizontal plane. This framework has been widely adopted for learning-based IO [8, 22] due to its effectiveness in simplifying pedestrian motion. Building on HACF, RIO [25] introduced rotation-equivalence data augmentation to self-supervise pose estimation after orientation alignment. Despite the success in pedestrian navigation and wheeled robots [9, 26], these representations are less effective for multirotor UAV motion, which require more sophisticated dynamic modeling. In this paper, we show that the commonly used global coordinate representations hinder neural networks from effectively capturing the dynamic complexities of UAVs.
II-C Inertial Odometry for Multirotor UAVs
Existing learning-based IO methods have primarily focused on pedestrian datasets, limiting their generalizability to platforms with more demanding motion dynamics, such as UAVs. Approaches like AI-driven pre-processing and down-sampling of high-speed inertial data have been proposed for better state estimation [27]. DIDO [15] estimates the thrust of UAVs using tachometer data and addresses unmodeled forces by incorporating a neural network. Building on this, the IMO [6] incorporates thrust information and IMU data within an EKF framework, which implicitly captures the drone dynamics. However, IMO relies on additional control signals to capture the drone’s dynamic, as shown in Fig. 8, which may suffer from overfitting on training datasets. DIVE [28], a more recent work for UAV IO, encodes orientation alongside gravity-removed global-frame acceleration. While this representation improves the robustness of learning-based IO, our findings in Section. V indicate that it remains suboptimal for capturing UAV dynamics comprehensively.
III Methodology
The goal of learning IO is to estimate the relative position transform (velocity) $\mathbf{v}$ from the IMU data $\{\mathbf{a}_{i},\mathbf{w}_{i}\}$ which is defined as $\mathbf{v}=\Psi(\{\mathbf{a}_{i},\mathbf{w}_{i}\})^{n}_{i=1}\in\mathbb{R}^{3}$ from the local body IMU frame. Most of the existing methods employ transforming the IMU frame to global frame or removing the gravity from the raw IMU data[9, 25, 28].
However, our study reveals that these widely used representations-such as global coordinate frames, HACF, and gravity-removed frames—limit the effectiveness of IMU feature extraction. Grounded on our analysis, our work consists of three parts: feature representation analysis, attitude encoding, and a tightly coupled EKF system.
We first investigate the feature representation of the learning IO in Section. III-A and Section. III-B, finding that the body coordinate frame is more representative than the other representation. Grounded on these findings, we propose AirIO which explicitly encode the drone’s pose to predict the body-frame velocity $\hat{\mathbf{v}}_{i}^{\mathcal{B}}$ and the corresponding uncertainty $\hat{\boldsymbol{\Sigma}}^{v}_{i}$ in Section. III-C. Lastly, in Section. III-D we build a tightly coupled EKF system that fuses the AirIO motion network with the learning-based IMU preintegration network, AirIMU [29], which jointly estimates the uncertainty of the IMU preintegration. Both modules predict their corresponding uncertainties, contributing to improved overall performance and robustness.
III-A IMU Coordinate Frame
Canonical Body Coordinate Frame: The IMU accelerometer measures the net force per unit mass acting on the sensor. Due to the measurement mechanism, it measures not only the object actual accelerations due to motion $\mathbf{\dot{v}}^{\mathcal{B}}=\frac{\mathbf{F}^{\mathcal{B}}_{{net}_{i}}}{m}$ in the sensor’s local coordinate frame (usually aligned with object body frame), but also the constant acceleration due to Earth’s gravity $\mathbf{g}^{\mathcal{G}}$ directed toward the Earth’s center. Thus, the accelerometer measurement at timestamp $i$ as:
$\mathbf{a}_{i}^{\mathcal{B}}=\frac{\mathbf{F}^{\mathcal{B}}_{{net}_{i}}}{m}+\mathbf{R}_{i}^{\top}\mathbf{g}^{\mathcal{G}},$ (1)
where $m$ is the object mass, and $\mathbf{R}_{i}\in\mathbb{SO}(3)$ is the object attitude, which defines the rotation for vector from body frame $\mathcal{B}$ to the global frame $\mathcal{G}$. $(\cdot)^{\mathcal{F}}$ denotes quantity represented in frame $\mathcal{F}$.
Figure 2: (a) Existing learning IO transforms the IMU input to the global coordinate using ground-truth orientation. (b) We found that preserving the IMU data in the body coordinate, along with gravitational information, can significantly improve accuracy. (c) Grounded on our previous analysis, our AirIO model leverages the body-frame IMU and explicitly encodes the drone’s orientation to estimate the body-frame velocity of the drone.
Global Coordinate Frame: Many existing methods transform the IMU measurement from body frame $\mathcal{B}$ to HACF or global coordinate frame $\mathcal{G}$ by applying an estimated transform rotation $\hat{\mathbf{R}}_{i}$, e.g.
$\mathbf{a}_{i}^{\mathcal{G}}=\hat{\mathbf{R}}_{i}\frac{\mathbf{F}^{\mathcal{B}}_{{net}_{i}}}{m}+\hat{\mathbf{R}}_{i}\mathbf{R}_{i}^{\top}\mathbf{g}^{\mathcal{G}}$ (2)
The rotation $\hat{\mathbf{R}}_{i}$ can often be obtained by simple IMU preintegration or the estimation results from an EKF, which is often accurate enough, such that $\hat{\mathbf{R}}_{i}\mathbf{R}_{i}^{\top}\approx\mathbf{I}$.
Notice that both (1) and (2) represents the same physical process. The key difference lies in how the attitude information is coupled with different components: in (1), it is coupled with the gravitational force, whereas in (2), it is coupled with the motion force. In (1), since the gravity acceleration $\mathbf{g}^{\mathcal{G}}$ is a constant, the the motion force and attitude form a linear combination. In contrast. in (2), the attitude couples with motion force in a nonlinear manner, with gravity acting as an additional constant that does not contribute extra useful information. Thus, the motion force and attitude are intuitively easier to estimate in (1), while (2) requires a more complex network structure for estimation.
III-B Representation Analysis
We perform feature analysis to verify our intuition. Specifically, we extract the latent features $\mathbf{h}_{i}$ from the IMU feature encoder for the input under different representations in Fig. 2. We conduct multiple feature analyses to assess their effectiveness and representativeness.
Principal Component Analysis PCA is a statistical method that reduces the dimensionality of a large data set by replacing correlated variables with a smaller set of uncorrelated principal components. To analyze the expressivity of the input representation, we performed PCA on the feature latent space for different input coordinates. We concatenate the latent features from all samples and form a feature matrix. Then, we perform PCA on the latent feature matrix to analyze its underlying structure by computing the singular value of the normalized feature matrix in a numerically stable way. Each singular value represents the corresponding magnitude scaling of one principal feature component. We then quantify the cumulative explained variance (or the energy coverage) for the top $k$ principal components in PCA. Specifically, we evaluate on two datasets. One is the widely used UAV dataset Blackbird[30] collected by a custom UAV platform Blackbird for agile autonomous flight, which covers large UAV maneuvers. The other is our custom simulation dataset Pegasus, which we collected using the NVIDIA IssicSim simulator with Pegasus autopilot [7] also with diverse maneuvers but in the ideal environment without external disturbance, allowing us to work with cleaner data.
Figure 3: The cumulative explained variance of the latent features for the Pegasus dataset (Left) and Blackbird dataset (Right). In both datasets, the latent features derived from body frames (black line) require fewer principal components to achieve comparable total energy.
Fig. 3 shows the cumulative energy coverage of the latent feature derived from two different IMU input representations: global and body frames. The red and black curves represent the energy coverage of the top $k$ principal components. Specially, for the Blackbird dataset, it shows that the top $k$ = 5 principal components in body-frame representation already cover 95% of the total energy. In contrast, global-frame representation requires the top $k$ = 8 principal components to achieve the same level of energy. Similarly, in the Pegasus datasets, the body-frame representation requires top $k$ = 15 principal components and the global frame require top $k$ = 40 principal components to cover 95% of the total energy. It clearly shows that models trained on the body representation input can achieve comparable performance with fewer features, highlighting the efficiency and expressivity of this representation in capturing the essential features. It also implies that models trained under body frame have better compressibility and can be designed more lightweight. Figure. 4 shows the comparison of the absolute trajectory error (ATE) for the models at different scales under both body and global frames. As the model size decreases, the prediction performance for model under body-frame representation degrades more smoothly and consistently yields lower errors, further showing the better compressibility and expressivity of the representation. More studies about the model compressibility can be found in Appendix. A-F.
Figure 4: The ATE for body-frame and global-frame model at different model sizes on Pegasus dataset (Left) and Blackbird dataset (Right). The text denotes the relative percentage increase in ATE.
t-SNE [31] We also perform t-SNE on Pegasus dataset, a dimensionality reduction approach that projects high-dimension data into a low-dimensional map to analyze the feature representation qualitatively. As shown in Fig. 5, based on the network prediction (here is the velocity), we color code each data point by the velocity magnitude from low to high, represented by blue to red. In the global coordinate frame, the points exhibit a highly entangled distribution and different velocity levels overlap, indicating poor separability and less representative encoding. In contrast, the feature distribution of the body coordinate frame is more structured and well-separated, where different velocity levels form more coherent clusters. This suggests that the body coordinate frame facilitates a more discriminative and effective representation of the latent features.
Figure 5: t-SNE analysis on Pegasus dataset with each point colored based on the velocity magnitude from low to high, represented by blue to red.
In summary, the analysis from Sec.III-A and Sec. III-B clearly indicates that i). A good prior information will help for estimation as it couples with the motion force in linear or nonlinear manner as shown in (1)(2). ii). The nonlinear coupling between the motion force and attitude in global-frame representation complicates the estimation problem. The body-frame representation proves to be more expressive and discriminative, leading to light network for prediction.
III-C AirIO Motion Network & Training
Based on the previous analysis, we introduce a separate encoding of the drone’s attitude to the network beyond the IMU encoder with body-frame representation, shown in Fig. 2(c). Inspired by [29], we map the drone’s orientation to the ${\mathfrak{so}(3)}$ Lie algebra space for compact, unconstrained, and continuous representation of 3D rotations and smoother network gradient to mitigate numerical instability.
We use CNN encoders to extract the features from the body-frame IMU data $(\mathbf{w}^{\mathcal{B}},\mathbf{a}^{\mathcal{B}})$ and the orientation of the drone $\boldsymbol{\xi}\in{\mathfrak{so}(3)}$. After concatenating the features from the attitude encoder and the IMU encoder, we employ bi-directional GRU layers to extract the latent feature $\mathbf{h}_{i}$ and model the temporal dependencies in the drone dynamics. In the decoder, we use two linear layers to output the body-frame velocity and its uncertainty. Overall the motion network can be written as: $(\hat{\mathbf{v}}^{\mathcal{B}},\hat{\boldsymbol{\sigma}}^{v})=f_{\boldsymbol{\theta}}(\mathbf{w}^{\mathcal{B}},\mathbf{a}^{\mathcal{B}},\boldsymbol{\xi})$. The model is supervised using an uncertainty loss inspired by [8]:
$L=(\mathbf{v}^{\mathcal{B}}_{i}-\hat{\mathbf{v}}^{\mathcal{B}}_{i})\hat{\boldsymbol{\Sigma}}^{v}_{i}{}^{-1}{(\mathbf{v}^{\mathcal{B}}_{i}-\hat{\mathbf{v}}^{\mathcal{B}}_{i}})^{T}+\ln(\det\hat{\boldsymbol{\Sigma}}^{v}_{i}),$ (3)
where $\hat{\boldsymbol{\Sigma}}^{v}_{i}$ are $3\times 3$ covariance matrix for the $i^{th}$ data. Similar to the setting in [8], we simplified the covariance as $\hat{\boldsymbol{\Sigma}}^{v}_{i}=\text{diag}(\hat{\boldsymbol{\sigma}}^{v}_{i}{}^{2})$, where $\hat{\boldsymbol{\sigma}}^{v}_{i}$ is learned uncertainty from the network. For orientation input $\boldsymbol{\xi}$, we use the ground truth orientations during training and replace them with EKF-estimated orientations during testing.
III-D Extended Kalman Filter
Figure 6: AirIO system pipeline. A tightly coupled EKF that fuses the AirIO motion network with learning-based IMU preintegration network AirIMU. Instead of using constant uncertainty parameters, we employ learned uncertainty $\hat{\boldsymbol{\sigma}}^{v}$ from the AirIO network and the learned gyroscope uncertainty $\hat{\boldsymbol{\sigma}}^{g}$ and accelerometer $\hat{\boldsymbol{\sigma}}^{a}$ in our EKF system.
Different from the existing IO methods [8, 6] that employ a motion network to capture the relative translation. In our implementation, our network models the body-frame velocity. This setup disentangles the current state of the drone from the previous state, which simplifies the EKF structure of our model. The full state of the filter is $\mathbf{X}=(\mathbf{R}^{\mathcal{G}}_{i},\mathbf{v}^{\mathcal{G}}_{i},\mathbf{p}^{\mathcal{G}}_{i},\mathbf{b}_{a_{i}},\mathbf{b}_{g_{i}})$, where $\mathbf{b}_{a_{i}}$ and $\mathbf{b}_{g_{i}}$ is the bias of the accelerometer and the gyroscope respectively. Following the common setup of the error-based filtering method, we define the error-state representation of the current filter state as $\delta\mathbf{X}=(\delta{\boldsymbol{\xi}}^{\mathcal{G}}_{i},\delta\mathbf{v}^{\mathcal{G}}_{i},\delta\mathbf{p}^{\mathcal{G}}_{i},\delta\mathbf{b}_{a_{i}},\delta\mathbf{b}_{g_{i}})$. For the rotation error, we define $\delta{\boldsymbol{\xi}}^{\mathcal{G}}_{i}=\mathbf{Log}(\mathbf{R}_{i}\hat{\mathbf{R}}_{i}^{-1})\in{\mathfrak{so}(3)}$, where $\mathbf{Log}(\cdot)$ denotes logarithm map operator.
AirIMU feature correction and uncertainty estimation To filter the noise and, more importantly, quantify the uncertainty of the IMU sensor model, we leverage AirIMU [29] to filter the raw IMU data. The model was trained through the differentiable IMU integrator and the covariance propagation to predict the corrections of the gyroscope and the accelerometer $[\boldsymbol{\sigma}_{g_{i}},\boldsymbol{\sigma}_{a_{i}}]\in\mathbb{R}^{3}$ with the corresponding uncertainty $[\boldsymbol{\eta}_{g_{i}},\boldsymbol{\eta}_{a_{i}}]\in\mathbb{R}^{3}$:
$\left(\hat{\boldsymbol{\sigma}}_{g_{i}},\hat{\boldsymbol{\sigma}}_{a_{i}}\right)=g_{\boldsymbol{\theta}}\left(\mathbf{w}_{i},\mathbf{a}_{i}\right),\ \ \ \left(\hat{\boldsymbol{\eta}}_{g_{i}},\hat{\boldsymbol{\eta}}_{a_{i}}\right)=\Sigma_{\boldsymbol{\theta}}\left(\mathbf{w}_{i},\mathbf{a}_{i}\right),$ (4)
We later filtered the IMU measurement by the predicted correction $\hat{\mathbf{a}}=\mathbf{a}+\hat{\boldsymbol{\sigma}}_{a_{i}},\hat{\mathbf{w}}=\mathbf{w}+\hat{\boldsymbol{\sigma}}_{g_{i}}$.
Filter Propagation The kinematic motion model of the filter propagation is:
$\begin{split}&\mathbf{R}_{i+1}=\mathbf{R}_{i}\cdot\mathbf{Exp}(\hat{\mathbf{w}}_{i}-\mathbf{b}_{g_{i}})\Delta t,\\ &\mathbf{v}_{i+1}=\mathbf{v}_{i}+\mathbf{R}_{i}(\hat{\mathbf{a}}_{i}-\mathbf{b}_{a_{i}}){\Delta}t,\\ &\mathbf{p}_{i+1}=\mathbf{p}_{i}+\mathbf{v}_{i}\Delta t+\frac{1}{2}\Delta t^{2}(\hat{\mathbf{a}}_{i}),\\ &\mathbf{b}_{a_{i+1}}=\mathbf{b}_{a_{i}},\mathbf{b}_{g_{i+1}}=\mathbf{b}_{g_{i}},\end{split}$ (5)
Here, the $\mathbf{Exp}(\cdot)$ denotes mapping from the log space to exponential space. The linearized propagation model is:
$\delta\mathbf{X}_{i+1}=\mathbf{A}_{i}\cdot\delta\mathbf{X}_{i}+\mathbf{B}_{i}\cdot\mathbf{n}_{i},$ (6)
where $\mathbf{n}_{i}=[\hat{\boldsymbol{\eta}}_{g_{i}},\hat{\boldsymbol{\eta}}_{a_{i}},\boldsymbol{\eta}_{b_{g}},\boldsymbol{\eta}_{b_{a}}]^{T}$. The $\hat{\boldsymbol{\eta}}_{g_{i}},\hat{\boldsymbol{\eta}}_{a_{i}}$ are the learned IMU uncertainty from the AirIMU network [29] with the $i$-th frame. The $\boldsymbol{\eta}_{b_{g}},\boldsymbol{\eta}_{b_{a}}$ are the random walk uncertainty set as a constant across different frames. The corresponding covariance propagation of the state covariance $\mathbf{P}_{i+1}$ follows:
$\mathbf{P}_{i+1}=\mathbf{A}_{i}\mathbf{P}_{i}\mathbf{A}_{i}+\mathbf{B}_{i}\mathbf{W}_{i}\mathbf{B}_{i}$ (7)
where the $\mathbf{W}_{i}$ is the covariance matrix of the $i$-th state, including the learned uncertainty of the IMU $\hat{\boldsymbol{\eta}}_{g_{i}},\hat{\boldsymbol{\eta}}_{a_{i}}$ and sensor bias random walk $\boldsymbol{\eta}_{b_{g}},\boldsymbol{\eta}_{b_{a}}$.
Filter Update The measurement update is as follows:
$h(\mathbf{X})=\mathbf{R}_{i}^{T}\cdot\mathbf{v}_{i}=\hat{\mathbf{v}}_{i}+\hat{\boldsymbol{\eta}}_{v_{i}}.$ (8)
where $\boldsymbol{\eta}_{v_{i}}$ is a random variable that follow the Gaussian distribution $\mathcal{N}(0,\hat{\boldsymbol{\Sigma}}^{v}_{i})$ learned by the network. The Jacobian matrix $\mathbf{H}$ is computed as:
$\begin{split}\mathbf{H}_{\delta v_{i}}&=\frac{\partial h(\mathbf{X})}{\partial\delta\mathbf{v}_{i}}=\mathbf{R}_{i}^{T}\\ \mathbf{H}_{\delta\xi_{i}}&=\frac{\partial h(\mathbf{X})}{\partial\delta\boldsymbol{\xi}_{i}}=\mathbf{R}_{i}^{T}[\mathbf{v}_{i}]_{\times}\end{split}$ (9)
Finally, the $\mathbf{H}$ and learned covariance are used in calculating the Kalman Gain:
$\begin{split}&\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{HPH}^{T}+\hat{\boldsymbol{\Sigma}}^{v}_{i})\\ &\mathbf{X}\leftarrow\mathbf{X}\oplus(\mathbf{K}(h(\mathbf{X})-\hat{\mathbf{v}}_{i}))\\ &\mathbf{P}\leftarrow(\mathbf{I}-\mathbf{KH})\mathbf{P}(\mathbf{I}-\mathbf{KH})^{T}+\mathbf{K}\hat{\boldsymbol{\Sigma}}^{v}_{i}\mathbf{K}^{T}\end{split}$ (10)
Operator $\oplus$ denotes the additional operation except for the orientation, where the update performs $\mathbf{R}\leftarrow\mathbf{Exp}(\boldsymbol{\xi})\cdot\mathbf{R}$.
IV Experiment
In this paper, we conduct evaluations on three quadrotor datasets: the real-world EuRoC dataset [32], the challenging drone racing dataset BlackBird [30], and our custom simulation dataset collected using Pegasus Simulator [7]. The Blackbird dataset is an aggressive indoor flight dataset. Compared to the EuRoC dataset, Blackbird presents more aggressive maneuvers and high-speed dynamics, making estimating the drone’s motion more challenging. The Pegasus dataset enables ablation studies with eliminating the ground-truth inaccuracies in real-world settings.
We compare AirIO with several methods. First, we select model-based IO methods like Baseline [14], the IMU preintegration approach in raw IMU data, and AirIMU [29], a hybrid method that corrects the raw IMU noise with data-driven method and integrates it with the IMU kinematic function. For learning-based IO, we compare against end-to-end trained model RoNIN [9] and an EKF-fused algorithm TLIO [8]. For recent learning-based IO methods targeting to UAVs, we compare against IMO [6], a control signal embedded IO specifically designed for drone racing. It incorporates additional thrust signals to capture drone motion.
IV-A Metrics Definition
We use the following metrics for evaluation:
Absolute Translation Error (ATE, $\mathrm{m}$)
$\text{ATE}=\sqrt{\frac{1}{N}\sum^{N}_{i=1}{\left\|\mathbf{p}_{i}-\hat{\mathbf{p}}_{i}\right\|_{2}^{2}}},$ (11)
calculates the average Root Mean Squared Error (RMSE) between the estimated and ground-truth positions over all time points.
Relative Translation Error (RTE, $\mathrm{m}$)
$\text{RTE}=\sqrt{\frac{1}{N}\sum^{N}_{i=1}\left\|\mathbf{p}_{i+\Delta t}-\mathbf{p}_{i}-\mathbf{R}_{i}\hat{\mathbf{R}}_{i}^{T}\left(\hat{\mathbf{p}}_{i+\Delta t}-\hat{\mathbf{p}}_{i}\right)\right\|^{2}_{2}},$ (12)
calculates the average RMSE of the relative displacements over predefined time intervals. Our evaluation specifically adopts a 5-second interval configuration.
IV-B Training Details
We use the Adam optimizer with an initialized learning rate of 0.001. The learning rate scheduler is implemented following the ReduceLROnPlateau, with patience of 5 epochs and a decay factor of 0.2. The batch size is 128. Our training and testing window size is set to 1000 frames (corresponding to 5 seconds in the EuRoC dataset). In the CNN encoder, we include a dropout layer with $p=0.5$ to reduce overfitting.
TABLE I: The ATE (Unit: ) and RTE (Unit: ) on the Blackbird dataset. Seen are sequences where the training and testing datasets are derived from the same trajectory. Unseen are those where the testing dataset includes sequences never used in training, demonstrating the model’s generalizability.
Seq. Baseline AirIMU RoNIN TLIO IMO AirIO Net AirIO EKF
ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE
Seen
    Clover 15.769 2.027 21.696 1.701 3.074 1.367 1.464 0.797 0.381 0.681 0.553 0.398 0.599 0.642
    Egg 66.170 7.677 13.675 4.178 2.449 2.552 2.227 2.398 1.153 0.828 0.649 0.310 0.549 0.307
    halfMoon 19.165 3.386 15.822 1.524 1.263 0.864 0.956 0.475 0.761 0.240 0.378 0.304 0.535 0.327
    Star 20.490 4.556 13.099 4.782 3.150 3.266 0.680 0.784 2.130 3.066 0.672 0.373 0.591 0.468
    Winter 15.781 2.395 11.800 1.772 1.031 1.089 0.616 0.755 0.219 0.206 0.396 0.183 0.332 0.196
Avg. 27.475 4.008 15.218 2.791 2.193 1.828 1.189 1.042 0.929 1.004 0.530 0.314 0.521 0.388
Unseen
    Ampersand 41.138 5.026 24.060 4.859 5.467 5.321 4.665 4.078 17.664 10.039 2.375 1.148 2.346 1.212
    Sid 13.108 2.055 16.601 2.782 18.581 10.665 7.323 8.427 9.967 6.992 0.683 0.539 0.686 0.531
    Oval 15.365 2.400 13.367 2.320 20.399 12.12 1.276 1.045 5.670 2.863 0.987 0.569 0.768 0.572
    Sphinx 3.429 2.713 10.737 3.736 11.537 7.762 2.005 2.408 5.629 5.395 1.204 1.020 1.119 1.017
    BentDice 28.307 2.342 38.060 3.036 11.453 6.792 2.119 1.747 6.143 4.519 1.156 0.955 1.111 1.226
Avg. 20.269 2.907 20.565 3.347 13.487 8.532 3.478 3.541 9.015 5.962 1.281 0.846 1.206 0.912
leverage ground truth orientation to transform the input data for inference.
Dead-reckoning IMU pre-integration.
IV-C System Performance
Blackbird Dataset Following IMO [6], we employ the same sequences from Blackbird and adopt their same train-test split configuration. We term these sequences as SEEN sequences because the training and testing data are from the same sequence. To evaluate generalization, we select five additional trajectories from BlackBird dataset as UNSEEN sequences, which are excluded from the training set. You may find more details on separation in Appendix. A-A1.
Figure 7: Trajectories of SEEN sequence halfMoon from the Blackbird dataset. AirIO, relying solely on IMU, outperforms IMO by 30% in ATE.
For the seen sequences, as shown in I, RoNIN and TLIO fail to capture the aggressive drone maneuvers as they are designed for pedestrian navigation with clear motion patterns. AirIO significantly outperforms these baselines, improving by 56.2% in ATE and 62.8% in RTE over TLIO. Furthermore, without integrating external thrust information like IMO, AirIO exceeds IMO by 43.9% in ATE and 61.4% in RTE relying solely on IMU measurement.
Surprisingly, we found that AirIO demonstrates remarkable generalizability to the unseen sequences. For example, as shown in 8, while other methods experience serious drift, AirIO maintains precise estimating. This performance gap arises because these methods tend to overfit training data, excelling within the training distribution but failing to generalize to UNSEEN data outside this distribution. Compared to IMO, AirIO shows a significant improvement of 86.6% in ATE and 84.7% in RTE.
Figure 8: Trajectories of UNSEEN sequence sid from the Blackbird dataset by RoNIN, TLIO, IMO, and AirIO.
EuRoC Dataset In Table. II, both RoNIN and TLIO show suboptimal performance compared to AirIO. After integrating EKF, AirIO further improves the accuracy, showing an average additional 17.4% improvement in ATE. Specifically, it achieves an improvement of 52.9% and 54.4% in ATE relative to RoNIN and TLIO, respectively. This success is largely attributed to effective uncertainty modeling. AirIO fully leverages the uncertainties learned from both the IMU kinematic model and the motion network, ensuring optimal data fusion and improved performance.
TABLE II: The ATE (Unit: ) and RTE (Unit: ) on EuRoC dataset.
Seq. RoNIN TLIO AirIO Net AirIO EKF
ATE RTE ATE RTE ATE RTE ATE RTE
MH02 5.902 2.121 7.281 2.451 4.917 0.936 2.478 0.987
MH04 8.586 4.542 8.626 5.498 2.726 1.093 2.308 1.005
V103 3.240 1.695 7.863 2.580 3.844 1.519 3.05 1.552
V202 7.445 2.303 6.260 2.783 4.823 1.303 4.206 1.313
V101 8.576 1.918 4.814 1.946 2.917 1.137 3.844 1.103
Avg. 6.75 2.516 6.969 3.052 3.846 1.198 3.177 1.192
leverage ground truth orientation to transform the input data for inference.
V Ablation Study
To evaluate the impact of different feature representations, we conduct ablation studies on the real-world dataset EuRoC and the simulation dataset Pegasus. We conduct experiments with the following input representation variants:
  1. i.
    Body: The network processes body-frame IMU measurements defined in Equation (1), as input signal.
  2. ii.
    Global: The network processes global-frame IMU as input. The global-frame IMU is transformed from raw IMU data, as defined in Equation (2).
  3. iii.
    Body + Attitude: Building upon i. Body, the network encodes drone orientation as auxiliary inputs.
  4. iv.
    Global + Attitude: Building upon ii. Global, the network encodes drone orientation as auxiliary inputs.
  5. v.
    Body – $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$: Building upon i. Body, we remove the rotation-coupled gravity term $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$ from Equation (1).
  6. vi.
    Global – $\mathbf{g}^{\mathcal{G}}$: Building upon ii. Global, we remove the gravity term $\mathbf{g}^{\mathcal{G}}$ from Equation (2).
To ensure a fair comparison, i. Body, iii. Body + Attitude, and v. Body – $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$ leverage ground-truth orientation to transform estimated velocity into global frame.
The results of the ablation study are presented in Table. III. You may find more results in Appendix. A-E. From these results, we address the following questions:
V-A How Effective is the Body-frame Representation?
Compared to ii. Global, i. Body demonstrates significantly higher precision on the Pegasus. This single step - switching from global frame to body frame - results in an average ATE improvement of 73.0%. i. Body on the real-world EuRoC dataset exhibits a similar trend, achieving a 53.1% improvement, as illustrated in Fig. 10.
We further evaluate ii. Global and i. Body using the Accuracy Area Under the Curve (AUC) metric, which quantifies the performance across various error thresholds. Specifically, we calculate the relative position error over a 5-second interval. As shown in 9, on both datasets, i. Body achieves higher accuracy across all error thresholds and reaches near-perfect accuracy much more quickly.
These results highlight how the body frame effectively captures implicit attitude information in IMU data. Simply encoding the body frame IMU features helps the network learn the drone’s dynamics. Moreover, we find that combining the body-frame representation with the attitude information iii. Body + Attitude is better than i. Body.
Figure 9: Accuracy AUC of EuRoC dataset(Left) and Pegasus dataset(Right).
Figure 10: Performance of MH_04_difficult in EuRoC dataset across different representations. ii. Global (Left) is suboptimal. Simply switching IMU data to i. Body (Center) improves significantly. After encoding drone rotation iii. Body + Attitude (Right), it has further improvement.
V-B How Effective is the Attitude Encoding?
Encoding the drone’s attitude information enhances the network’s ability to capture the dynamics of drones. As illustrated in Fig. 10, on the EuRoC dataset, iii. Body + Attitude achieves a further 30.3% improvement than the i. Body, outperforming ii. Global by 67.3% in ATE.
Additionally, we evaluate attitude encoding on global-frame representation iv. Global + Attitude. This also yields an improvement of approximately 29.6% over ii. Global, validating the efficiency of encoding attitude information.
V-C Shall We Keep Gravity in the Feature Representation?
Many learning-based IO methods remove gravity from the IMU encoding [28]. However, as discussed in Section. III-A, gravitational acceleration is a critical component for body-frame representation as it explicitly embeds drone rotation through the term $\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$, as shown in Equation 1.
After removing this rotation-coupled gravity term, i. Body will discard essential attitude information. Table. III demonstrates that v. Body –$\mathbf{R}^{T}_{i}\mathbf{g}^{\mathcal{G}}$ causes significant performance degradation: ATE increases by 90.0% on EuRoC and 98.7% on Pegasus dataset.
TABLE III: Ablation study on the EuRoC and Pegasus datasets comparing different feature representations. Evaluation metric: ATE (Unit: ).
Seq. Global (2) $-\mathbf{g}^{\mathcal{G}}$ Global Body (1) $-\mathbf{R}_{i}^{T}\mathbf{g}^{\mathcal{G}}$ Global $+$Attitude Body Body $+$Attitude
EuRoC
MH02 18.048 17.892 16.136 9.522 4.537 2.531
MH04 15.904 6.895 8.977 4.572 3.561 2.250
V103 7.007 7.086 6.586 4.068 6.824 3.107
V202 9.662 3.496 6.161 10.274 3.248 4.310
V101 8.142 15.11 7.082 7.117 5.478 4.287
Avg. 11.753 10.096 8.988 7.110 4.730 3.297
Pegasus
TEST_1 22.115 9.689 9.534 6.151 5.167 3.418
TEST_2 15.745 15.185 14.729 5.373 4.862 5.719
TEST_3 19.467 26.960 3.580 14.545 3.982 1.786
Avg. 19.109 17.278 9.281 8.690 4.670 3.641
VI Conclusion & Discussion
In this work, we address the challenges identified in the discussion of IMO[6] without extra information like thrust signal. We demonstrate that simply preserving body-frame IMU data, inclusive of gravitational acceleration, significantly improves IO performance on UAVs. Building upon this analysis, we propose the AirIO system, which further enhances accuracy by explicitly encoding attitude information and integrating it with an EKF. Notably, compared to IMO, our model generalizes effectively to trajectories not present in the training datasets and operates without relying on additional inputs such as thrust. We believe that the proposed work has broader implications for reliable state estimation in agile drone flight.
Appendix A Appendix
A-A BlackBird Dataset
The Blackbird unmanned aerial vehicle (UAV) dataset is an indoor dataset[30] designed to capture aggressive flight maneuvers for fully autonomous drone racing. It collects data using a Blackbird quadrotor platform with an Xsens MTi-3 IMU. The blackbird takes place in a motion capture room and follows a predefined periodic trajectory, each lasting approximately 3-4 minutes at high speed.
A-A1 Seen and Unseen sequences separation
[…]
A-A2 Training and Testing Sequences Separation
In our experiments, we follow the same dataset-splitting strategy presented in IMO: for each trajectory, the data is allocated as 70% for training, 15% for validation, and the remaining 15% for testing. We use the SEEN sequences’ training and validation set to train our model, making our training setup identical to IMO’s. For testing, we employ both the SEEN sequences’ testing set and the UNSEEN sequences’ testing set. The comprehensive testing setup allows us to evaluate our method’s generalization to new trajectories.
TABLE IV: Separation of trajectory sequences into SEEN and UNSEEN categories, and their respective allocations to training, validation, and testing sets.
SEEN clover Egg halfMoon Star Winter
training (70%)
validation (15%)
testing (15%)
UNSEEN Ampersand Sid Oval Sphinx BentDice
training (70%)
validation (15%)
testing (15%)
A-B Qualitative Evaluation for Blackbird dataset
We present more details on the evaluation of the Blackbird dataset. As shown in 11, we showcase seven additional trajectories that further highlight our method’s performance. Our method achieves superior performance in the SEEN sequences. For more than half of these sequences, it outperforms existing methods that rely on additional information in addition to IMU measurements, while our method uses only IMU data. When evaluating UNSEEN sequences, our model outperforms all existing methods on all sequences, demonstrating its remarkable adaptability.
(a) SEEN Sequences. From top to bottom: Clover, Egg, Winter and Star sequence. Our method demonstrates robust performance without requiring any additional sensor information.
(b) UNSEEN Sequences. From top to bottom: bentDice, Oval, and Sphinx sequence. Our method demonstrates remarkable adaptability to trajectories it has never seen before.
Figure 11: Estimated trajectories of Blackbird dataset by RoNIN, TLIO, IMO and AirIO (Ours).
A-C Pegasus Dataset
We collected a simulation dataset in the open-source Pegasus Simulator [7] to evaluate our proposed method under controlled conditions. Pegasus is a framework built on top of NVIDIA Omniverse and Isaac Sim. It is designed for multirotor simulation and supports integration with PX4 firmware, as well as Python control interfaces. In our setup, we used QGroundControl to control the multirotor and also employed the quadratic thrust curve and linear drag model, ensuring the generated flight dynamics closely resemble real-world conditions.
In our experiment, we collected a total of seven trajectories datasets, named Pegasus Dataset. We divided the Pegasus dataset into training and testing sets. Specifically, four trajectories are selected for training and the remaining three trajectories are testing. They are illustrated in 12 and 13. We provide a detailed overview of each trajectory as follows.
[…]
Figure 12: Training set
[…]
Figure 13: Testing set
A-D EuRoC Dataset
The EuRoC datasets are the well-known benchmarks for odometry and SLAM algorithms. They are collected by a micro aerial vehicle: an AscTec Firefly hex-rotor helicopter. There are 11 trajectories collected in two scenarios: an industrial environment and a motion capture room. We selected MH_01_easy, MH_03_medium, MH_05_difficult, V1_02_medium, V2_01_easy, V2_03_difficult for training, and the rest for testing.
Figure 14: The MH_04_difficult trajectories from the EuRoC dataset visualized within its 3D reconstruction map. While RoNIN (dark green) and TLIO (light green) fail, AirIO (red) retains a coherent trajectory shape.
A-E Ablation Study in Pegasus and EuRoC dataset
TABLE V: Ablation study on the EuRoC and Pegasus datasets comparing different feature representations. Evaluation metric: RTE (Unit: ).
Seq. Global -- gravity Global Body -- gravity Global $+$Attitude Body Body $+$Attitude
EuRoC
MH02 1.684 1.575 2.346 1.542 1.314 0.972
MH04 2.618 1.961 2.525 1.707 1.329 1.009
V103 1.407 1.352 1.613 1.485 1.623 1.512
V202 1.721 1.789 2.176 1.723 1.373 1.263
V101 1.801 1.991 1.463 1.498 1.122 1.104
Avg. 1.846 1.734 2.025 1.591 1.352 1.172
Pegasus
TEST_1 2.783 2.971 2.134 1.694 1.561 1.017
TEST_2 2.704 3.007 1.961 2.339 2.314 1.905
TEST_3 3.274 3.350 1.298 1.969 0.672 0.396
Avg. 2.920 3.109 1.798 2.001 1.516 1.106
A-F Ablation Study on Model Compression
To evaluate the compressibility of different representations, we introduced additional two lightweight models Light A and Light B. The light models keep the same layer structure but shrink the dimensions of each layer’s hidden units. Finally, the encoder’s latent feature dimension is reduced from 256 to 128, and then further to 64, yielding progressively smaller models.
To quantify the performance degradation as the model is compressed, we define the degradation ratio for ATE and RTE. A higher degradation ratio indicates a larger drop in performance. As shown in VI, the model under body frame shows smoother degradation in both ATE and RTE.
TABLE VI: Ablation study on the Blackbird, Pegasus, and EuROC datasets comparing compressibility of models under body frame and global frame. Evaluation metric: ATE (Unit: ), RTE(Unit: ), and Degradation.
Model Regular Light A Light B
Feature Size 256$\times$1 128$\times$1 64$\times$1
Model Size 2.524 MB 0.641 MB 0.175 MB
Metrics ATE Degradation ATE Degradation ATE Degradation
Blackbird Body 0.647 - 0.755 16.8% 0.931 44.0%
Global 0.952 - 1.238 29.9% 1.522 59.8%
Pegasus Body 4.670 - 10.118 116.6% 15.192 225.3%
Global 17.278 - 30.950 79.1% 69.236 300.7%
EuRoC Body 4.730 - 5.447 15.2% 6.875 45.4%
Global 10.096 - 14.033 39.0% 38.236 278.7%
Metrics RTE Degradation RTE Degradation RTE Degradation
Blackbird Body 0.345 - 0.454 31.3% 0.510 47.7%
Global 0.544 - 0.583 7.2% 0.983 80.6%
Pegasus Body 1.516 - 2.226 46.9% 2.203 45.3%
Global 3.109 - 3.422 10.0% 4.858 56.2%
EuRoC Body 1.352 - 1.359 0.5% 1.297 -4.1%
Global 1.734 - 2.176 25.5% 4.468 157.8%